OpenStudy (anonymous):

Write the standard form of the line that contains a slope of 2/3 and passes through the point (1, 1).

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

okay so you know m=2/3 \[y-y _{1}=m(x-x _{1})\] can you do this and we will work from there?

OpenStudy (anonymous):

???

OpenStudy (amorfide):

let's try a different approach you have m=2/3 m is the slope btw and you know the y value from the coordinates given y=1 you have a line \[y=\frac{ 2 }{3 }x + b\]

OpenStudy (amorfide):

\[1=\frac{ 2 }{3 }x + b\]

OpenStudy (amorfide):

you know x=1 too sub that in and work out b

OpenStudy (amorfide):

\[1=\frac{ 2 }{ 3 }(1) + b\]

OpenStudy (amorfide):

is that more clearer?

OpenStudy (anonymous):

So how do I put it into standard form?